9. (1/5 Points] DETAILS PREVIOUS ANSWERS LARLINALG8 7.1.025. Find the characteristic equation and the eigenvalues and correspplz solve all 3
Answer #1

Solution: (9) Given : A = -3 -4 4 3 6 4 The characteristic equation | A-Il = 0 oma -H 子。子。 -3 3 4-7 -6 4-) -3 3 4-) -6 4-2 -4- 6 (42) (1-62 -9A +2y) = 0 (2+ 2) (2 ( 3-6 ) –ฯ (3-4)) ( 4+ ) (3-6) (3-4) - 6 -V - 0 A = 9 (6) Eigenvalues (4,2 , 003) I (-2si X = . 3/22] 312) le 23 la dz V = 3/2 I X, na 4 -4 -4 -3 3 - 6 R₂-R, R2 TO or é o wie oor mio moro mro .4 -3 3 Re R2 Too I3 6 -4 o -3 -2 -6 o -2 2 -3 11 R₂-4R, R2 (R2 - 2 R, > R2) -2 R3-R2 R3 3 -8 4 Re -8 is o R2 = 3 R2-Rio Ri [: оор > Ri = 1 / sl10 (10) A - १ म JA-AT | = 0 4 3 - 2 - O ५- (5-2) (3-2)(५->)-0 +4 (0) = (s-2) (3-2) (4-2) --0 5--0 3--0 ५- ८० a=3 -५ Eigenvalu